OpenStudy (anonymous):

use separation of variables to solve the initial value problem dy/dx=5-x^2/3y; y=-1 when x=0

OpenStudy (anonymous):

\[3yy'=5-x^2\] \[3\int \frac{y'}{y}dx=\int 5-x^2 dx\\ 3\ln(y)=5x-\frac{x^3}{3}\] is a start

OpenStudy (anonymous):

i did \[3ydy=5-x^2dx\] \[\int\limits3ydy=\int\limits5-x^2dx \] \[3y^2/2=5x-x^3/3+c\]